# Adaptive Mix for Semi-Supervised Medical Image Segmentation

### v1

@@ 3.1 Problem statement
Before delving into the proposed method, we first provide the notations that will be used subsequently. The training set $\mathcal{D}=\{\mathcal{D^{L}},\mathcal{D^{U}}\}$ contains a labeled set $\mathcal{D^{L}}=\{(x^{l}_{i},y^{l}_{i})_{i=1}^{N^{L}}\}$ and an unlabeled set $\mathcal{D^{U}}=\{(x^{u}_{j})_{j=1}^{N^{U}}\}$, where $x^{l}_{i}$/$x^{u}_{j}$ denotes the $i_{th}$/$j_{th}$ labeled/unlabeled image, $y^{l}_{i}$ is the ground truth of the labeled image, and $N^{L}$ and $N^{U}$ ($N^{U}>>N^{L}$) are the numbers of labeled and unlabeled samples. Given the training data $\mathcal{D}$, semi-supervised semantic segmentation aims to learn a model $f(\cdot;\theta)$ performing well on unseen test sets $\mathcal{T}$.
Figure 2: Overview of the proposed Adaptive Mix framework for semi-supervised medical image segmentation. It includes our Adaptive Mix (AdaMix) algorithm for generating perturbed images and a semi-supervised learning paradigm for providing pseudo labels. AdaMix can be seamlessly applied to the self-training, mean-teacher, and co-training paradigms, named AdaMix-ST, AdaMix-MT, and AdaMix-CT, respectively.
@@ 3.2 Overview
As discussed in Section 1, the random region mix-up (Yun et al., 2019) does not allow for adaptively controlling perturbation degree, leading to the learning model overfitting to randomly perturbed examples. Meanwhile, the confidence-guided mix (Shen et al., 2023), which utilizes a fixed mix-up rule that replaces low-confidence patches with high-confidence ones, produces mixed images with trivial perturbation degrees, thereby limiting the effect of consistency regularization. In contrast, replacing high-confidence regions with low-confidence regions may generate new samples with overly strong perturbations, making it difficult for the model to adapt to these new samples. We pinpoint the question: How can image mix-up augmentation be adaptively performed during training? To this end, we propose Adaptive Mix (AdaMix), a novel augmentation algorithm for adaptive image mix-up based on a model’s learning state. Building upon AdaMix, we construct an Adaptive Mix framework, where AdaMix acts as a plug-in-play module and can be seamlessly integrated into the self-training, mean-teacher, and co-training paradigms for semi-supervised medical image segmentation. In the following, we first specify the Self-Pacecd Mix framework [Fig. 2] and then collaborate on the Self-Pacecd Mix algorithm [Fig. 3].
@@
Output: $f(\cdot;\theta)$
@@
2:     # Forward propagation (w/o gradient) to obtain pseudo labels for unlabeled original/auxiliary images $x^{u}_{j}/x^{u}_{jj}$
3:     $\tilde{y}^{u}_{j}\leftarrow f(x^{u}_{j},\theta)$ and $\tilde{y}^{u}_{jj}\leftarrow f(x^{u}_{jj},\theta)$
@@
5:     $\bar{x}^{l}_{i},\bar{y}^{l}_{i}\leftarrow\mathrm{AdaMix}(x^{l}_{i},x^{l}_{ii},y^{l}_{i},y^{l}_{ii},f(\cdot,\theta))$
6:     $\bar{x}^{u}_{j},\bar{y}^{u}_{j}\leftarrow\mathrm{AdaMix}(x^{u}_{j},x^{u}_{jj},\tilde{y}^{u}_{i},\tilde{y}^{u}_{ii},f(\cdot,\theta))$
7:     # Forward propagation to obtain predictions for labeled and unlabeled perturbed images $\bar{x}^{l}_{i}$ and $\bar{x}^{u}_{j}$
8:     $\hat{y}^{l}_{i}\leftarrow f(\bar{x}^{l}_{i},\theta)$ and $\hat{y}^{u}_{j}\leftarrow f(\bar{x}^{u}_{j},\theta)$
9:     # Loss supervision by Eq. 1
10:     Calculate $\mathcal{L}_{s}$ based on $\hat{y}^{l}_{i}$, $\bar{y}^{l}_{i}$
11:     Calculate $\mathcal{L}_{u}$ based on $\hat{y}^{u}_{j}$, $\bar{y}^{u}_{j}$
@@ 3.3 Adaptive Mix Framework
As illustrated in Fig. 2, the AdaMix framework includes two components: 1) the AdaMix algorithm for generating perturbed examples, and a semi-supervised learning scheme for conducting supervision on both labeled and unlabeled images. According to the manners of strong-weak pseudo supervision (Sohn et al., 2020), semi-supervised learning can be realized through three paradigms: 1) self-training, 2) mean-teacher, and 3) co-training, in which FixMatch (Sohn et al., 2020), Mean-Teacher (Tarvainen and Valpola, 2017), and CPS (Chen et al., 2021) are the corresponding representation frameworks, respectively.
@@
Formally, these paradigms can be formulated as follows. Given $(x^{l}_{i},y^{l}_{i})\in\mathcal{D^{L}}$, $x^{u}_{j}\in\mathcal{D^{U}}$ and its corresponding pseudo label $\tilde{y}^{u}_{j}$ with the confidence map $\tilde{p}^{u}_{j}$, the semi-supervised learning objective can be defined as:
$J=\mathop{\min}\limits_{\theta}\underbrace{\mathcal{L}_{seg}(f(x^{l}_{i},\theta),y^{l}_{i})}_{\mathcal{L}_{s}}+\underbrace{\mathbb{I}_{\{\tilde{p}^{u}_{j}\geq\tau\}}(\tilde{p}^{u}_{j})\mathcal{L}_{seg}(f(\Gamma(x^{u}_{i}),\theta),\tilde{y}^{u}_{j})}_{\mathcal{L}_{u}}\$ (1)
@@
where the first and second terms denote the supervised loss $\mathcal{L}_{s}$ and the unsupervised loss $\mathcal{L}_{u}$, respectively. $\mathcal{L}_{seg}(\cdot,\cdot)$ is an image segmentation criterion, $f(\cdot,\theta)$ denotes a segmentation model with learnable parameters $\theta$, $\tau$ refers to a confidence threshold above which a pixel-level pseudo-label is retained for pseudo supervision. $\Gamma(\cdot)$ represents a strong perturbation/augmentation, which is the key to the efficacy of consistency regularization.
@@
Self-Training involves a single model $f(\cdot,\theta)$, which generates pseudo labels from unlabeled data for training of the model itself. A pseudo label is generated by: $\tilde{y}^{u}_{j}=\arg\max{\mathrm{SoftMax}(f(\Gamma^{\prime}(x^{u}_{i}),\theta))}$, where $\mathrm{SoftMax}(\cdot)$ denotes the SoftMax function that converts predictive logits into probabilities and $\Gamma^{\prime}(\cdot)$ represents a weak perturbation (w.r.t. the strong perturbation $\Gamma(\cdot)$)  [^2] Following the idea of strong-weak pseudo supervision (Sohn et al., 2020), in practice, $\Gamma^{\prime}(\cdot)$ refers to ’weak’ perturbations, generally consisting of only geometric augmentations, while $\Gamma(\cdot)$ refers to ’strong’ augmentations, which include geometric and pixel degree augmentations, as well as mix-up operations..
Figure 3: Schematic diagram of the proposed Adaptive Mix algorithm (AdaMix). It performs image mix-up perturbation in a self-paced manner, synthesizing samples from easy to hard during training based on the model’s learning state.
@@
Mean-Teacher includes a student model $f(\cdot,\theta)$ and a teacher $f(\cdot,\tilde{\theta})$ model, where the parameters $\tilde{\theta}$ of the teacher model is updated by an exponential moving average (EMA) of the student’s parameters $\theta$ in each iteration $t$: $\tilde{\theta}_{t}=\alpha\tilde{\theta}_{t-1}+(1-\alpha)\theta_{t}$. Pseudo labels are generated by the teacher model: $\tilde{y}^{u}_{j})=\arg\max{\mathrm{SoftMax}(f(\Gamma^{\prime}(x^{u}_{i}),\tilde{\theta}))}$.
@@
Co-Training involves two (student) models, $f(\cdot,\theta_{1})$ and $f(\cdot,\theta_{2})$, which are cross-supervised from each other on unlabeled data. The pseudo labels yielded by one model are leveraged to supervise another model, i.e., $\tilde{y}^{u}_{j[1/2]})=\arg\max{\mathrm{SoftMax}(f(\Gamma^{\prime}(x^{u}_{i}),\theta_{[1/2]}))}$.
@@
Building upon these paradigms and the strong-weak pseudo supervision, we develop three frameworks with our AdaMix, i.e., AdaMix-ST, AdaMix-MT, and AdaMix-CT, for semi-supervised medical image segmentation. These instantiated frameworks employ Eq. 1 as their learning objective and exploit AdaMix as the strong perturbation $\Gamma(\cdot)$ for consistency regularization. The training procedure of these methods is detailed in Algorithm 1 (Note that, without loss of generality, we take AdaMix-ST as an example for illustration convenience). We will show in Section 4 that the proposed frameworks archive state-of-the-art performance on various medical image segmentation tasks.
@@
Input: Original image $x^{u}_{j},\tilde{y}^{u}_{j}$ and auxiliary image $x^{u}_{jj},\tilde{y}^{u}_{jj}$
Output: AdaMix image $\bar{x}^{u}_{j},\bar{y}^{u}_{j}$
@@
2:  $\hat{y}^{u}_{j},z^{u}_{j}\leftarrow f(x^{u}_{j},\theta)$ and $\hat{y}^{u}_{jj},z^{u}_{jj}\leftarrow f(x^{u}_{jj},\theta)$
@@
4:  Calculate $\bar{\mathcal{L}}$ based on $\hat{y}^{u}_{j},\bar{y}^{u}_{j}$ and $\hat{y}^{u}_{jj},\bar{y}^{u}_{jj}$
@@
6:  Update $m$ by Eq. 4 to determine the Mix-Up rule $L\leftarrow H$ or $H\leftarrow L$
@@
8:  Update $v$ by Eq. 6 to determine the number of Mix-Up patches $n$
@@
11:     $\bar{x}^{u}_{j}/\bar{y}^{u}_{j}\leftarrow$ Replace top $n$ low-confidence patches in $x^{u}_{j}/\bar{y}^{u}_{j}$ with top $n$ high-confidence patches in $x^{u}_{jj}/\bar{y}^{u}_{jj}$
@@
13:     $\bar{x}^{u}_{j}/\bar{y}^{u}_{j}\leftarrow$ Replace top $n$ high-confidence patches in $x^{u}_{j}$ with top $n$ low-confidence patches in $x^{u}_{jj}/\bar{y}^{u}_{jj}$
@@ 3.4 Adaptive Mix algorithm
Given an unlabeled original image $x^{u}_{j}$, an unlabeled auxiliary image $x^{u}_{jj}$, and their corresponding pseudo labels $\tilde{y}^{u}_{j},\tilde{y}^{u}_{jj}$, we first obtain their confidence maps by $z^{u}_{j}=\max(\mathrm{SoftMax}(f(x^{u}_{j},\theta)))$ and $z^{u}_{jj}=\max(\mathrm{SoftMax}(f(x^{u}_{jj},\theta)))$, respectively. Meanwhile, we calculate the self-paced proxy loss ¯=¯$\bar{\mathcal{L}}=\bar{\mathcal{L}}_{o}+\bar{\mathcal{L}}_{a}$, where ¯$\bar{\mathcal{L}}_{o}=\mathcal{L}_{seg}(f(x^{u}_{j},\theta),\tilde{y}^{u}_{j})$ and ¯$\bar{\mathcal{L}}_{a}=\mathcal{L}_{seg}(f(x^{u}_{jj},\theta),\tilde{y}^{u}_{jj})$ denote the segmentation loss of the original image-pseudo label pair and the auxiliary image-pseudo label pair, respectively. Based on the self-paced proxy loss $\bar{\mathcal{L}}$, we can define the self-paced mask and self-paced weight as follows.
@@
where $g(v,\lambda)=\lambda(\frac{1}{2}v^{2}-v)$ is a linear self-paced regularizer. By solving Eq. 5, we obtain the solution for $v$:
@@
Finally, based on the confidence maps $z^{u}_{j}$ and $z^{u}_{jj}$, the mix-up rule $r$, and the number of mix-up patches $n$, we perform patch mix-up with patch size $S$ to derive an AdaMix perturbed example $(\bar{x}^{u}_{j},\bar{y}^{u}_{j})$. Please refer to Algorithm 2 for a detailed description of the procedure above.
4 Experiments and Results
4.1 Datasets
4.2 Implementation Details
Framework: We employ U-Net (Ronneberger et al., 2015)/V-Net (Milletari et al., 2016) as the supervised baseline architecture for 2D/3D image segmentation, respectively. We utilize FixMatch (Sohn et al., 2020), MT (Tarvainen and Valpola, 2017), and CPS (Chen et al., 2021) as the base frameworks for self-training, mean-teacher, and co-training, respectively. These frameworks build upon the concept of strong-weak pseudo supervision (Sohn et al., 2020) to leverage unlabeled data. We set the confidence threshold $\tau=0.95$ following FixMatch (Sohn et al., 2020). For the hyperparameters of AdaMix, we set the mix-up patch size $S=32$ and the maximum number of mix-up patches $K=16$. Detailed analysis for these hyperparameters is provided in Section 4.5. In our experiments, we adopt a combination of cross-entropy and dice loss (Milletari et al., 2016) as the segmentation criterion.
Table 1: Comparison with state-of-the-art methods under the 10% $|\mathcal{D}|$ and 20% $|\mathcal{D}|$ labeled data on the ACDC dataset. Note that all methods adopt U-Net as the backbone. The best and second-best results are highlighted in red and blue, respectively.
@@ 4.2 Implementation Details
CPS (Chen et al., 2021) 85.84 76.15 7.06 2.17 87.07 78.11 5.38 1.82 CutMix Co-Training
@@
Table 2: Comparison with state-of-the-art methods on the LA dataset with 10% $|\mathcal{D}|$ and 20% $|\mathcal{D}|$ labeled data. Note that all methods adopt V-Net as the backbone. The best and second-best results are highlighted in red and blue, respectively.

### v2

@@ 3.1 Problem statement
Before delving into the proposed method, we first provide the notations that will be used subsequently. The training set $\mathcal{D}=\{\mathcal{D}^{L},\mathcal{D}^{U}\}$ contains a labeled set $\mathcal{D}^{L}=\{(x^{l}_{i},y^{l}_{i})_{i=1}^{N^{L}}\}$ and an unlabeled set $\mathcal{D}^{U}=\{(x^{u}_{j})_{j=1}^{N^{U}}\}$, where $x^{l}_{i}$/$x^{u}_{j}$ denotes the $i^{th}$/$j^{th}$ labeled/unlabeled image, $y^{l}_{i}$ is the ground truth for the labeled image, and $N^{L}$ and $N^{U}$ ($N^{U}>>N^{L}$) are the numbers of labeled and unlabeled samples. Given the training data $\mathcal{D}$, semi-supervised semantic segmentation aims to learn a model $f(\cdot;\theta)$ that performs well on unseen test sets $\mathcal{T}$.
@@ 3.2 Overview
As discussed in Section 1, the random region mix-up (Yun et al., 2019) does not allow for adaptively controlling perturbation degree, leading to the learned model overfitting to randomly perturbed examples. Meanwhile, the confidence-guided mix (Shen et al., 2023), which utilizes a fixed mix-up rule that replaces low-confidence patches with high-confidence ones, produces mixed images with trivial perturbation degrees, thereby limiting the effect of consistency regularization. In contrast, replacing high-confidence regions with low-confidence regions may generate new samples with overly strong perturbations, making it difficult for the model to adapt to these samples. We pinpoint the question: How can image mix-up augmentation be adaptively performed during training? To this end, we propose Adaptive Mix (AdaMix), a novel augmentation algorithm for adaptive image mix-up based on a model’s learning state. Building upon AdaMix, we devise an Adaptive Mix framework, where AdaMix acts as a plug-and-play module and can be seamlessly integrated into the self-training, mean-teacher, and co-training paradigms for semi-supervised medical image segmentation. In the following, we first specify the Adaptive Mix framework [Fig. 2] and then elaborate on the Adaptive Mix algorithm [Fig. 3].
@@
Output: Trained segmentation model $f(\cdot;\theta)$
@@
2:     # Forward propagation (w/o gradient) to obtain pseudo labels for unlabeled original/auxiliary images $x^{u}_{j}/x^{u}_{b}$ ($j\neq b$)
3:     $\bar{y}^{u}_{j}\leftarrow f(x^{u}_{j},\theta)$ and $\bar{y}^{u}_{b}\leftarrow f(x^{u}_{b},\theta)$
@@
5:     $\tilde{x}^{l}_{i},\tilde{y}^{l}_{i}\leftarrow\mathrm{AdaMix}(x^{l}_{i},x^{l}_{a},y^{l}_{i},y^{l}_{a},f(\cdot,\theta))$
6:     $\tilde{x}^{u}_{j},\tilde{y}^{u}_{j}\leftarrow\mathrm{AdaMix}(x^{u}_{j},x^{u}_{b},\bar{y}^{u}_{j},\bar{y}^{u}_{a},f(\cdot,\theta))$
7:     # Forward propagation to obtain predictions for labeled and unlabeled perturbed images $\tilde{x}^{l}_{i}$ and $\tilde{x}^{u}_{j}$
8:     $\hat{y}^{l}_{i}\leftarrow f(\tilde{x}^{l}_{i},\theta)$ and $\hat{y}^{u}_{j}\leftarrow f(\tilde{x}^{u}_{j},\theta)$
9:     # Loss supervision by Eq. (1)
10:     Calculate $\mathcal{L}_{s}$ based on $\hat{y}^{l}_{i}$, $\tilde{y}^{l}_{i}$
11:     Calculate $\mathcal{L}_{u}$ based on $\hat{y}^{u}_{j}$, $\tilde{y}^{u}_{j}$
@@
Figure 3: Schematic diagram of the proposed Adaptive Mix algorithm (AdaMix). It performs image mix-up perturbation in a self-paced manner, synthesizing perturbed samples from easy to hard during training based on the model’s learning state.
@@ 3.3 Adaptive Mix Framework
As illustrated in Fig. 2, the AdaMix framework includes two components: 1) the AdaMix algorithm for generating perturbed examples and 2) a semi-supervised learning scheme for conducting supervision on both labeled and unlabeled images. According to the approach of strong-weak pseudo supervision (Sohn et al., 2020), semi-supervised learning can be realized through three paradigms: 1) self-training, 2) mean-teacher, and 3) co-training, in which FixMatch (Sohn et al., 2020), Mean-Teacher (Tarvainen and Valpola, 2017), and CPS (Chen et al., 2021b) are the corresponding representative frameworks, respectively.
@@
Formally, these paradigms can be formulated as follows. Given $(x^{l}_{i},y^{l}_{i})\in\mathcal{D}^{L}$, $x^{u}_{j}\in\mathcal{D}^{U}$ and its corresponding pseudo label $\bar{y}^{u}_{j}$ with the confidence map $\bar{p}^{u}_{j}$, the semi-supervised learning objective can be defined as:
$J=\mathop{\min}\limits_{\theta}\mathbb{E}_{\mathcal{D}}[\underbrace{\mathcal{L}_{seg}(f(x^{l}_{i},\theta),y^{l}_{i})}_{\mathcal{L}_{s}}+\underbrace{\mathbb{I}_{\{\bar{p}^{u}_{j}\geq\tau\}}(\bar{p}^{u}_{j})\mathcal{L}_{seg}(f(\Gamma(x^{u}_{j}),\theta),\bar{y}^{u}_{j})}_{\mathcal{L}_{u}}]\$ (1)
@@
where the first and second terms denote the supervised loss $\mathcal{L}_{s}$ and the unsupervised loss $\mathcal{L}_{u}$, respectively. $\mathcal{L}_{seg}(\cdot,\cdot)$ is an image segmentation criterion, $f(\cdot,\theta)$ denotes a segmentation model with learnable parameters $\theta$, and $\tau$ refers to a confidence threshold above which a pixel-level pseudo label is retained for pseudo supervision. $\Gamma(\cdot)$ represents a strong perturbation/augmentation, which is the key to the efficacy of consistency regularization.
@@
Self-Training involves a single model $f(\cdot,\theta)$, which generates pseudo labels from unlabeled data for training of the model itself. A pseudo label is generated by: $\bar{y}^{u}_{j}=\arg\max{\mathrm{SoftMax}(f(\Gamma^{\prime}(x^{u}_{j}),\theta))}$, where $\mathrm{SoftMax}(\cdot)$ denotes the SoftMax function that converts predictive logits into probabilities and $\Gamma^{\prime}(\cdot)$ represents a weak perturbation (w.r.t. the strong perturbation $\Gamma(\cdot)$)  [^2] Following the idea of strong-weak pseudo supervision (Sohn et al., 2020), in practice, $\Gamma^{\prime}(\cdot)$ refers to ”weak” perturbations, generally consisting of only geometric augmentations, while $\Gamma(\cdot)$ refers to ”strong” augmentations, which include geometric and pixel intensity augmentations, as well as mix-up operations. For brevity, these notations are omitted in the following text..
@@
Mean-Teacher includes a student model $f(\cdot,\theta)$ and a teacher $f(\cdot,\bar{\theta})$ model, where the parameters $\bar{\theta}$ of the teacher model is updated by an exponential moving average (EMA) of the student’s parameters $\theta$ at each iteration $t$: $\bar{\theta}_{t}=\alpha\bar{\theta}_{t-1}+(1-\alpha)\theta_{t}$. Pseudo labels are generated by the teacher model: $\bar{y}^{u}_{j}=\arg\max{\mathrm{SoftMax}(f(\Gamma^{\prime}(x^{u}_{j}),\bar{\theta}))}$.
@@
Co-Training involves two (student) models, $f(\cdot,\theta_{1})$ and $f(\cdot,\theta_{2})$, which are cross-supervised from each other on unlabeled data. The pseudo labels yielded by one model are leveraged to supervise another model, i.e., $\bar{y}^{u}_{j[1/2]}=\arg\max{\mathrm{SoftMax}(f(\Gamma^{\prime}(x^{u}_{j}),\theta_{[1/2]}))}$.
@@
Building upon these paradigms and the strong-weak pseudo supervision, we develop three frameworks with our AdaMix, i.e., AdaMix-ST, AdaMix-MT, and AdaMix-CT, for semi-supervised medical image segmentation. These instantiated frameworks employ Eq. (1) as their learning objective and exploit AdaMix as the strong perturbation $\Gamma(\cdot)$ for consistency regularization. The training procedure of these methods is detailed in Algorithm 1 (Note that, without loss of generality, we take AdaMix-ST as an example for illustration convenience). We will show in Section 4 that the proposed frameworks achieve state-of-the-art performance on various medical image segmentation tasks.
@@
Input: Original image $(x^{u}_{j},\bar{y}^{u}_{j})$ and auxiliary image $(x^{u}_{b},\bar{y}^{u}_{b})$ ($j\neq b$)
Output: AdaMix image $\tilde{x}^{u}_{j},\tilde{y}^{u}_{j}$
@@
2:  $\hat{y}^{u}_{j},\hat{p}^{u}_{j}\leftarrow f(x^{u}_{j},\theta)$ and $\hat{y}^{u}_{b},\hat{p}^{u}_{b}\leftarrow f(x^{u}_{b},\theta)$
@@
4:  Calculate $\bar{\mathcal{L}}$ based on $\hat{y}^{u}_{j},\bar{y}^{u}_{j}$ and $\hat{y}^{u}_{b},\bar{y}^{u}_{b}$
@@
6:  Update $m$ by Eq. (4) to determine the Mix-Up rule $L\leftarrow H$ or $H\leftarrow L$
@@
8:  Update $v$ by Eq. (6) to determine the number of Mix-Up patches $n$
@@
11:     $\tilde{x}^{u}_{j}/\tilde{y}^{u}_{j}\leftarrow$ Replace top $n$ low-confidence patches in $x^{u}_{j}/\bar{y}^{u}_{j}$ with top $n$ high-confidence patches in $x^{u}_{b}/\bar{y}^{u}_{b}$
@@
13:     $\tilde{x}^{u}_{j}/\tilde{y}^{u}_{j}\leftarrow$ Replace top $n$ high-confidence patches in $x^{u}_{j}/\bar{y}^{u}_{j}$ with top $n$ low-confidence patches in $x^{u}_{b}/\bar{y}^{u}_{b}$
@@ 3.4 Adaptive Mix algorithm
Given an unlabeled original image $x^{u}_{j}$, an unlabeled auxiliary image $x^{u}_{b}$, and their corresponding pseudo labels $\bar{y}^{u}_{j},\bar{y}^{u}_{b}$, we first obtain their confidence maps by $\hat{p}^{u}_{j}=\max(\mathrm{SoftMax}(f(x^{u}_{j},\theta)))$ and $\hat{p}^{u}_{b}=\max(\mathrm{SoftMax}(f(x^{u}_{b},\theta)))$, respectively. Meanwhile, we calculate the self-paced proxy loss ¯=¯$\bar{\mathcal{L}}=\bar{\mathcal{L}}_{o}+\bar{\mathcal{L}}_{a}$ as an indicator of the model’s state, where ¯$\bar{\mathcal{L}}_{o}=\mathcal{L}_{seg}(f(x^{u}_{j},\theta),\bar{y}^{u}_{j})$ and ¯$\bar{\mathcal{L}}_{a}=\mathcal{L}_{seg}(f(x^{u}_{b},\theta),\bar{y}^{u}_{b})$ denote the segmentation loss for the original image-pseudo label pair and the auxiliary image-pseudo label pair, respectively. Based on the self-paced proxy loss $\bar{\mathcal{L}}$, we define the self-paced mask and self-paced weight as follows.
@@
where $g(v,\lambda)=\lambda(\frac{1}{2}v^{2}-v)$ is a linear self-paced regularizer. By solving Eq. (5), we obtain the solution for $v$:
@@
Finally, based on the confidence maps $\hat{p}^{u}_{j}$ and $\hat{p}^{u}_{b}$, the mix-up rule $r$, and the number of mix-up patches $n$, we perform patch mix-up with patch size $S$ to derive an AdaMix perturbed example $(\tilde{x}^{u}_{j},\tilde{y}^{u}_{j})$. Please refer to Algorithm 2 for a detailed description of the procedure above.
Table 1: Comparison with state-of-the-art methods under the 10% $|\mathcal{D}|$ and 20% $|\mathcal{D}|$ labeled data on the ACDC dataset. Note that all methods adopt U-Net as the backbone. The best and second-best results are highlighted in red and blue, respectively. indicates statistically significant improvements ($p$-value $<0.05$ based on a paired t-test) of our AdaMix frameworks over ABD, MT, and UCMT, respectively.
@@
CPS (Chen et al., 2021b) 85.84 76.15 7.06 2.17 87.07 78.11 5.38 1.82 CutMix Co-Training
@@
4 Experiments and Results
4.1 Datasets
4.2 Implementation Details
Framework: We employ U-Net (Ronneberger et al., 2015)/V-Net (Milletari et al., 2016) as the supervised baseline architecture for 2D/3D image segmentation, respectively. We utilize FixMatch (Sohn et al., 2020), MT (Tarvainen and Valpola, 2017), and CPS (Chen et al., 2021b) as the base frameworks for self-training, mean-teacher, and co-training, respectively. These frameworks are built upon the concept of strong-weak pseudo supervision (Sohn et al., 2020) to leverage unlabeled data. We set the confidence threshold $\tau=0.95$ following FixMatch (Sohn et al., 2020). For the hyperparameters of AdaMix, we set the mix-up patch size $S=32$ and the maximum number of mix-up patches $K=16$. Detailed analysis for these hyperparameters is provided in Section 4.4.3. In our experiments, we adopt the dice loss (Milletari et al., 2016) as the segmentation criterion.
Data: In the 2D image segmentation tasks, all images are resized to $256\times 256$ for inference, while the outputs are recovered to their original sizes for evaluation. For 3D segmentation, we randomly crop $80\times 112\times 112\,(Depth\times Height\times Width)$ patches for training and iteratively crop patches using a sliding window strategy to obtain the final segmentation mask for evaluation.
4.3 Comparison with State of the Arts
We extensively compared our proposed AdaMix frameworks with state-of-the-art semi-supervised learning methods, which can be divided into four groups based on their perturbation strategies: 1) Multi-task perturbation (using signed distance maps (Li et al., 2020; Luo et al., 2021a) or multi-scale features (Luo et al., 2021b)): SASSNet (Li et al., 2020), URPC (Luo et al., 2021b), and DTC (Luo et al., 2021a); 2) Noise perturbation: UA-MT (Yu et al., 2019), MC-Net (Wu et al., 2021), MC-Net+ (Wu et al., 2022); 3) CutMix perturbation: FixMatch (Sohn et al., 2020), MT (Tarvainen and Valpola, 2017), CPS (Chen et al., 2021b); and 4) UMix and I-UMix perturbations: UCMT (Shen et al., 2023) and ABD (Chi et al., 2024).
Table 2: Comparison with state-of-the-art methods on the LA dataset with 10% $|\mathcal{D}|$ and 20% $|\mathcal{D}|$ labeled data. Note that all methods adopt V-Net as the backbone. The best and second-best results are highlighted in red and blue, respectively. indicates statistically significant improvements ($p$-value $<0.05$ based on a paired t-test) of our AdaMix frameworks over ABD, MT, and UCMT, respectively.